INAF queue observing, 2018 December 14-15

Observer: R. Carini, A. Rossi <br />
Telescope Operator: S. Allanson <br />
Support Astronomer: B. Rothberg <br />
Instrument: LBC

Summary

Night duration (12\xB0 el twilight) 1:15-13:15 UT = 12h 00m

Observing time: 8h 00min (63%)

Weather time loss: 4h 0m (33%)

technical time loss 0h 30m (4%)

Data Summary

At the sunset the sky is completely covered.

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images Notes
CALIBRATIONS arcflats MODS1,2 mask 31 dual grating     Gilli mask 31
telluric Nisini field 4 HIP16652 luci1,2 MOS G210 HiRes H J K (lambda 2.2) K (2.22)   luci1.0017-0032, luci2.0020-0035  
Nisini Field 4 luci1, 2 MOS G210 HiRes H J K 2h/3h luci1.0040-0055, luci2.0042-0057 we take more images due to clouds, completed
telluric field2 HIP17971 Luci1,2 MOS G210 HiRes H J K (lambda 2.2) K (2.22)   luci1.0059-0074, luci2-0061-0076  
Nisini Field 2 luci1, 2 MOS G210 HiRes K, J 1h/3h luci1.0081-0097, luci2.0085-0100 passing cirrus, to be checked
telluric Marziani HIP52192 luci1, 2       useless
telluric Nardini q6 luci1, 2     luci1.0102-0107 passing cirrus
Nardini q6 luci1 150 K 0.25/0.67 luci1.0113-0116 passing cirrus check please
CALIBRATIONS   Nisini field 4, 2, 3 arcflats

Details

00:00 UT Overcast,we stay close

MODS CALIBRATIONS
  • Gilli mask31 dualgrating arcflats mods1r.20181215.0002-0010,mods1b.20181215.0002-0010; mods2r.20181215.0002-0010,mods2b.20181215.0002-0010

03:14 the sky is a little bit clearer, we open.

03:31 send the script for the telluric HIP16652, Nisini field4
  • acq luci1.2011215.0001-0002, luci1.20181215.0004-0005
  • the exposure time is too low for this sky covered. we don't see the telluric in the image. We take another image a time exposure of 2.5(dit)x10(ndit)
  • acq luci1.0003-0005, luci2,0006-0008
  • 03:48 error copoiniting limit, right side.
  • 03:50 we sand again the offset calculated during the alignment
  • sci ==luci.0006-0013, luci2.0008-0016
  • there is something wrong int he script, we have no spectra in the rightest slit.

*04:17 send again the acquisition
  • abort and check pointing
  • 04:20 send again the script luci1.0014-0016, luci2.0017-0019
  • during the acquisition luci2 didn't take the image of the star, we take it manually.
  • sci luci1.0017-0032, luci2.0020-0035

05:10 Slew to Nisini Field 4
  • seeing 1" on guider but 2 mag of extinction, estimated from the guiding star
  • wait for target to transit at culmination, otherwise is >88 degrees
  • 05:20 send again script
  • acq luci1.0033-0039, luci2.0036-0041
  • 05:23 problems with guinding try again
  • 05:37 we cannot acquire there are problems: sky iamge is missing and MOS allignement gives errors.
  • 05:38 send script again
  • seeing 1 hole in the sky almost clear at 5:45UT
  • 05:54 start sci luci1.0040-0044, luci2.0042-0045 K filter==
  • some cirrus but dimm seeing 1.0"
  • luc1.0045-0049, luci2.0046-0050 J filter (we take one more image in both side because for clouds passed during that integration)
  • luci1.0050-0055, luci2.0051-0057

07:05 slew to telluric HIP17971 for the field 2
  • acq luci1.0056-0058, luci2.0058-0060
  • sci luci1.0059-0074, luci2-0061-0076
  • 07:20 error grating, we follow the instructions in the troubleshooting page
  • 07:we go ahead with script even if the centrale wavelenght in luci2 is not stable. NOw it works.
  • clouds are coming, 3 mag of extintion

08:00 slew to Nisini field2 seeing 1.1-1.4, passing cirrus
  • acq luci1.0075-0080, luci2.0077-0084
  • the maskhas severasl probelms: 1) one allignement slit 33 is not in the right position. Must have been edited manually; 2) we had to add -0.2(luci1) and -0.5(luci2)degrees to center the mask and several offsets. (negative rotation means moving stars counterclockwise)
  • 08:35 sent sci luci1.0081-0097, luci2.0085-0100
  • seeing 1.2 on guider 08:36, passing cirrus
  • lost guiding at08:50 we pause
  • 08:56 we continue
  • We did only K and J bands

10:03 We slew to telluric HIP52192.xml for Marziani
  • check pointing
  • 10:13 mask in fpu
  • problem moving mask
  • seeing , passing dense cirrus

10:40 Nardini q6 get visible. Move to telluric HIP 53606
  • 10:45 we cannot guide , too cloudy, we wait.
  • acq luci1.0100-0101
  • sci luci1.0102-0107 passing cirrus, difficult to guide , seeing ~1
  • we took 2 images more because of the low signal, but the trace is obvious.

11:09 Slew to Nardini q6
  • acq luci1.0109-0111
  • seeing 0.8
  • sci luci1.0112 guide lost, too cloudy , not good
  • 11:25 we wait
  • 12:00 guide is back we try again
  • sci luci1.0113-0116 wesee the trace, 117-118 are not good

12:15 overcast, thick clouds, TO closes the dome and take calibrations.

CALIBRATIONS
  • Nisini field2 K, 2.2 luci1.00120-0135, luci2.0101-0116
  • Nisini field2 K, 2.22 luci1.0136-0151, luci2.0117-0132
  • Nisini field3 K, 2.22 luci1.0152-167, luci2.0133-0148
  • Nisini field4 H luci1.0168-182, luci2.0149-0162
  • Nisini field4 J luci1.0183-195, luci2.0163-0176
  • Nisini field4 K2.20 luci1.0196-0211, luci2.0177-0192
  • Nisini field4 K2.22 luci1.0212-02, luci2.0193-01
Topic revision: r11 - 20 Dec 2018, AndreaRossi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback